I can't figure out how to set this problem up. "One type of antifreeze is 40% gylcol and another type of antifreeze is 60% gylcol. How much of each kind should be used to make 100 gallons of antifreeze that is 48% glycol." Help please. Thanks.
Answer by Paul(988)

40x+60(100-x)=48(100)
40x+6000-60x=4800
-20x=-1200
x=60

100-60=40

hence, for 40% about 60 G is used and for 60% about 40 G is used.
Paul.